Circular No. 7740

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fb IN ANONYMOUS GALAXY A. Rest, University of Washington, reports the detection of an apparent

supernova (at magnitudes g′ = 20.0, r′ = 19.0) on Sloan Digital Sky Survey (SDSS) images taken with the SDSS 2.5-m reflector at Apache Point on Oct. 10.3 UT. SN 2001fb is located at α = 0^{h}10^{m}06^{s}.15, δ = −0^{o}26′18′′.2 (equinox 2000.0), which is 7′′.2 west and 8′′.7 south of the nucleus of the host galaxy. The new object is not present in SDSS images taken on 1998 Sept. 19.3, to limiting magnitudes g′ = 23 and r′ = 22.5. Follow-up observations taken at the ARC 3.5-m reflector at Apache Point Observatory on Oct. 24.1 by A. Miceli and R. Covarrubias confirm the existence of the new object.

SUPERNOVA 1998ff IN ANONYMOUS GALAXY S. Beckwith, Space Telescope Science Institute, on behalf of the ‘Super-

novae at High Z Team’ (which includes also A. Riess, F. Boffi, S. Casertano, R. Doxsey, H. Ferguson, A. Fruchter, M. Giavalisco, R. Gilliland, I. Griffin, A. Koekemoer, M. Livio, B. Margon, G. Meylan, N. Panagia, V. Platais, K. Sahu, and D. Soderblom), reports the discovery of an apparent supernova in the ‘Hubble Deep Field South’ (HDFS) on WFPC2 images taken with the Hubble Space Telescope (HST) and filter F814W in Sept. and Oct. 1998: “SN 1998ff was found by differencing the 1998 images with new HST images taken during 2001 Oct. 2 and 3, also with F814W. The new object appeared with a starlike point-spread function, offset 0′′.08 west of the core of an early-type galaxy (number 2285 from Casertano et al. 2000, A.J. 120, 2747) whose nucleus is located at α = 22^{h}32^{m}48^{s}.90, δ = −60^{o}34′04′′.8 (equinox 2000.0; uncertainty ± 0′′.04). SN 1998ff had a mean magnitude of [F814W] = 25.3 ± 0.1 during the HDFS campaign and was seen to decline by a few tenths of a magnitude during the 12 days of the campaign. SN 1998ff was independently detected in each of eight separate, 1-day epochs between 1998 Sept. 28 and Oct. 9. A photometric redshift of z = 1.20 for the host galaxy was given by Fontana et al. (2000, A.J. 120, 2206) by employing photometry from ultraviolet wavelengths through the K band. Low-resolution spectroscopy of the host galaxy (so as to refine further the host-galaxy redshift and knowledge of the underlying stellar population), as well as analysis of deep images of the HDFS taken in the later half of 1998, may prove valuable.”

2001 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green